# Subgroups of Cyclically Amalgamated Free Products

### v1

@@ 1. Introduction
Amalgamated free products arise naturally in group theory and other branches of mathematics. For instance, surface groups (i.e., fundamental groups of compact 2-dimensional manifolds), various knot groups, and the well-known Baumslag-Solitar groups are of this type. For two factors, amalgamated free products were first introduced in 1927 by O. Schreier [Sch]. For several factors, this was later generalized by H. Neumann [Neu1, Neu2]. W. Magnus was one of the first mathematicians who recognized their value and used them in the early 1930s in several papers on one-relator groups [Mag1, Mag2].
It is an important task to describe the subgroups of amalgamated free products and the relation between their structure and properties of the group itself. A first major step in this direction was taken by A. Karrass and D. Solitar in [KS], where the authors give a very general and rather abstract classification which is, however, not easy to apply in concrete situations. In this paper we examine the structure of subgroups with a small number of generators and the existence of free subgroups and free products of cyclic groups in the following type of group.
@@
In [FRW], the authors showed that, for $n\in\{2,3\}$, every $n$-generated subgroup in certain groups of F-type is a free product of cyclics. Later it was discovered that this classification can be suitably extended to 4-generated subgroups. In Section 3 we generalize these results as follows (see Proposition 10 and Theorems 11 and 12).
@@
Here we say that a group is an $n$-free product of cyclics if every $n$-generated subgroup of $G$ is a free product of at most $n$ cyclic groups. Notice that this notion is entirely different from the the classical notion of $n$-free groups, as introduced by G. Higman in 1951 (cf. [Hig]). In particular, the factors of a product of at most $n$ cyclic groups are allowed to be finite.
Moreover, we point out that case (a) of this theorem has been known for a long time (see [KS2, Thm. 6]) and is included here merely for completeness sake.
@@
Next we discuss low rank subgroups of cyclically amalgamated free products in Section 3. Besides proving the above theorem, we also apply it to recover the known results about low rank subgroups of groups of F-type (see Corollary 13). In the proof of these results, the method of Nielsen reductions plays a central role, as it allows us to keep a tight control over the structure of the words representing the group elements under consideration. It has been suggested that, possibly, the method of groups acting on trees, as developed by H. Bass and J.P. Serre in [BS], and in particular in the form elaborated by R. Weidmann (cf. [Wei]), could be used to obtain a proof of the above theorem. Except for case (a) of the theorem, which has been previously known, this is highly doubtful, as in this case the Bass-Serre method encounters certain obstructions which are hard to analyze in detail (see [Wei, Thm. 2]) and do not lead to explicit presentations of the subgroups under consideration in a conceivable way.
@@
The central conjecture here is the Surface Group Conjecture which asks whether an infinite, finitely presented, non-free group, such that every subgroup of infinite index is free, has to be isomorphic to a surface group. A major advance in this area was recently achieved by H. Wilton who proved the conjecture (among others) for one-relator groups using heavy topological machinery [Wil2]. In Theorem 16, we point out a proof in the case of cyclically amalgamated free products which relies chiefly on group theoretic methods.
@@
Throughout we have strived to keep this paper as self-contained as possible. All definitions and statements used without reference can be found in [FMRSW].
@@ 2. Nielsen Reduction in Amalgamated Free Products
In this section we describe the Nielsen reduction method in amalgamated free products. This method was originally developed by H. Zieschang in [Zie1] and then refined by the third author in [Ros1, Ros2] and, together with R.N. Kalia, in [KR].
@@
Next we fix a system $L_{i}$ of left coset representatives of $A$ in $H_{i}$ for $i=1,2$. Here we represent $A$ by $1$. The following proposition is well-known (see [FMRSW, Thm. 1.5.8 and Cor. 1.5.10]).
@@
For $g\in G$ with reduced form $g=h_{1}\cdots h_{m}a$, the number $m$ is called the length of $G$. For our purposes, the partial ordering for $G$ given by the length is too coarse. Thus we have to refine it. For $i=1,2$, we take the elements of $L_{i}^{-1}$ as a system of representatives for the right cosets of $A$ in $H_{i}$. Now we have the following normal form (see [FMRSW, Sect. 1.5.2] or [ZVC, Prop. 2.3.12]).
@@
    The set $\{g_{1},\dots,g_{n}\}$ is called shorter than the set $\{h_{1},\dots,h_{n}\}$ if $g^{\prime}_{i}\preccurlyeq h^{\prime}_{i}$ for $i=1,\dots,n$ and if $h^{\prime}_{j}\prec g^{\prime}_{j}$ fails for at least one $j\in\{1,\dots,n\}$.
@@
Using these rules, the following result was shown by H. Zieschang in [Zie1, Satz 1 and Korollar 1] and refined by the third author in [Ros1, Ros2].
@@
Finally, we recall some refinements of the preceding material which were shown in [KR] and will be useful later. For an element $g\in G$ with symmetric normal form $g=\ell_{1}\cdots\ell_{m}\,k\,r_{m}\cdots r_{1}$ as in Proposition 3, we write $\ell(g)=\ell_{1}\cdots\ell_{m}$ for the leading half, $k(g)=k$ for its kernel, and $r(g)=r_{m}\cdots r_{1}$ for its rear half. Thus we have $g=\ell(g)\cdot k(g)\cdot r(g)$.
@@
Claim (a) is shown in [KR, Lemma 2.4] and claim (b) in [KR, Lemma 2.5].
@@ 3. Low Rank Subgroups of Amalgamated Free Products
Notice that this notion is completely different from the concept of an “$n$-free group” which usually means that every $n$-generated subgroup is free (see [Hig]).
In the following we examine subgroups of an amalgamated free product $G=H_{1}\underset{A}{\ast}H_{2}$ which satisfies Assumption 2. In particular, let $H_{1},H_{2}$ be finitely generated groups, and let $A=H_{1}\cap H_{2}$ be an infinite cyclic group which is malnormal in $G$. In this setting, 2-generated subgroups of $G$ can be characterized explicitly. Note that the following result was first shown in [KS2, Thm. 6] and is included here for completeness sake.
@@
Parts (a) and (b) of this corollary were previously shown in [FRW]. The next example shows that part (c) is optimal.
@@
Let $G$ be a group of F-type which is not a free product of cyclics. Assume that every subgroup of $G$ of infinite index is a free product of cyclics. Then $G$ is a co-compact planar discontinuous group (cf. [FMRSW, Thm. 3.5.45]).
@@ 4. Free Subgroups of Infinite Index
The central open question for free subgroups of infinite index is the Surface Group Conjecture. The first version of this conjecture was formulated 1980 in the Kourovka Notebook by O.V. Melnikov (cf. [KM, Problem 7.36]). It has since been disproved, but numerous improved versions were formulated over time (for instance, see [FKMRR, CFR, FRW]), and the latest one reads as follows (cf. [Wil2]).
@@
A celebrated result of H. Wilton is that the surface group conjecture holds for infinite one-relator groups (cf. [Wil2]). Its proof is based on a number of deep results in topology. Another important situation where this conjecture holds is the case of fully residually free groups (see [CFR]).
@@
For this kind of group, the surface group conjecture was shown in [CFR, Thm. 3.1.1]. For the convenience of the reader, we briefly recall the main steps. We distinguish four cases depending on whether $u$ and $v$ are proper powers of elements of $G$.
@@
Case 4: Assume that neither $u$ nor $v$ is a proper power in $G$. In this case, [JR, Thm. A] shows that $G$ is a word hyperbolic group. Next we use [Sta, Thm. 4.11] to show that $G$ is one-ended. In fact, if $G$ has no ends, then $G$ is finite. If $G$ has two ends, then $G$ has an infinite cyclic subgroup of finite index. Finally, if $G$ has more than two ends, it has infinitely many ends and is a non-trivial free product. As all of these cases are impossible for the groups $G$ under consideration, they have to be one-ended. The final step is an application of H. Wilton’s result that a hyperbolic, one-ended cyclically pinched one-relator group, all of whose subgroups of infinite index are free, is a surface group (cf. [Wil1, Corollary 4]).
@@
Not much seems to be known for low rank subgroups and free subgroups of infinite index of HNN extensions. For 2-generated and 3-generated subgroups, the following result from [FRR] is the best we know. If we drop the assumption that $u$ is not conjugate in $H$ to $v$ or $v^{-1}$, a slightly stronger version for 2-generated subgroups of $G$ is given in [FMRSW, Thm. 1.6.42].
@@
As regards to the surface group conjecture, we can show it for the following type of HNN extensions. Notice that the following result is slightly stronger than [CFR, Thm. 3.1.2] and [FMRSW, Thm. 1.7.22].
@@
Let $N$ be the normal closure of $H$ in $G$. Then $G/N$ is an infinite cyclic group. Hence the group $H$ is of infinite index in $G$, and the hypothesis implies that $H$ is a free group. Thus $G$ is a conjugacy pinched one-relator group. For such groups, the surface group conjecture can be shown based in part on [CFR, Thm. 3.1.2]. Let us follow the main steps and complete the proof at the appropriate place. We distinguish two cases depending on whether $u$ or $v$ are proper powers of elements of $G$.
Case 1: Assume that $u=g^{m}$ and $v=h^{n}$ with $g,h\in G$ and with $m,n\geq 2$. If $G=\langle g,h\mid g^{2}=h^{2}\rangle$ then $G$ is a surface group and contains a free abelian subgroup of rank 2. Now assume that $G$ is not isomorphic to $\langle g,h\mid g^{2}=h^{2}\rangle$. Using the normal form in HNN groups (cf. [FMRSW, Thm. 1.6.4]), it follows that the subgroup generated by $w=tgt^{-1}$ and $h$ has the presentation $W=\langle w,h\mid w^{m}=h^{n}\rangle$. Then the subgroup $\langle w^{m},wh\rangle$ of $W$ is a free abelian group of rank 2 and has infinite index in $G$, in contradiction to the hypothesis.
@@
In the case $\mathop{\rm rk}\nolimits(H)=1$, we let $H=\langle a\rangle$. Then we get $G=\langle a,t\mid tat^{-1}=a^{k}\rangle=B(1,k)$ with a Baumslag-Solitar group $B(1,k)$. For $|k|=1$, this is a surface group, and for $p=|k|\geq 2$, the group $B(1,p)$ is isomorphic to $\mathbb{Z}[\frac{1}{p}]\rtimes\mathbb{Z}$ (cf. [FM]) and contains $\mathbb{Z}[\frac{1}{p}]$ as a non-free subgroup of infinite index. Hence, also in the case $G\cong B(1,k)$ with $|k|\geq 2$, the group $G$ has a non-free subgroup of infinite index, in contradiction to the hypothesis.
Subcase 2b: Assume that $\langle u\rangle\cap x\langle v\rangle x^{-1}$ is finite for every $x\in H$. Since $G$ is torsionfree, we get $\langle u\rangle\cap x\langle v\rangle x^{-1}=\{1\}$. By [KMy], it follows that $G$ is word hyperbolic. As in the proof of Theorem 16, we see that $G$ is one-ended, and then an application of [Wil1] finishes the proof.

### v2

@@ 1. Introduction
Amalgamated free products arise naturally in group theory and other branches of mathematics. For instance, surface groups (i.e., fundamental groups of compact 2-dimensional manifolds), various knot groups, and the well-known Baumslag-Solitar groups are of this type. For two factors, amalgamated free products were first introduced in 1927 by O. Schreier [22]. For several factors, this was later generalized by H. Neumann [17, 18]. W. Magnus was one of the first mathematicians who recognized their value and used them in the early 1930s in several papers on one-relator groups [15, 16].
It is an important task to describe the subgroups of amalgamated free products and the relation between their structure and properties of the group itself. A first major step in this direction was taken by A. Karrass and D. Solitar in [11], where the authors give a very general and rather abstract classification which is, however, not easy to apply in concrete situations. In this paper we examine the structure of subgroups with a small number of generators and the existence of free subgroups and free products of cyclic groups in the following type of group.
@@
In [7], the authors showed that, for $n\in\{2,3\}$, every $n$-generated subgroup in certain groups of F-type is a free product of cyclics. Later it was discovered that this classification can be suitably extended to 4-generated subgroups. In Section 3 we generalize these results as follows (see Proposition 10 and Theorems 11 and 12).
@@
Here we say that a group is an $n$-free product of cyclics if every $n$-generated subgroup of $G$ is a free product of at most $n$ cyclic groups. Notice that this notion is entirely different from the the classical notion of $n$-free groups, as introduced by G. Higman in 1951 (cf. [8]). In particular, the factors of a product of at most $n$ cyclic groups are allowed to be finite.
Moreover, we point out that case (a) of this theorem has been known for a long time (see [12, Thm. 6]) and is included here merely for completeness sake.
@@
Next we discuss low rank subgroups of cyclically amalgamated free products in Section 3. Besides proving the above theorem, we also apply it to recover the known results about low rank subgroups of groups of F-type (see Corollary 13). In the proof of these results, the method of Nielsen reductions plays a central role, as it allows us to keep a tight control over the structure of the words representing the group elements under consideration. It has been suggested that, possibly, the method of groups acting on trees, as developed by H. Bass and J.P. Serre in [23], and in particular in the form elaborated by R. Weidmann (cf. [25]), could be used to obtain a proof of the above theorem. Except for case (a) of the theorem, which has been previously known, this is highly doubtful, as in this case the Bass-Serre method encounters certain obstructions which are hard to analyze in detail (see [25, Thm. 2]) and do not lead to explicit presentations of the subgroups under consideration in a conceivable way.
@@
The central conjecture here is the Surface Group Conjecture which asks whether an infinite, finitely presented, non-free group, such that every subgroup of infinite index is free, has to be isomorphic to a surface group. A major advance in this area was recently achieved by H. Wilton who proved the conjecture (among others) for one-relator groups using heavy topological machinery [26]. In Theorem 16, we point out a proof in the case of cyclically amalgamated free products which relies chiefly on group theoretic methods.
@@
Throughout we have strived to keep this paper as self-contained as possible. All definitions and statements used without reference can be found in [5].
@@ 2. Nielsen Reduction in Amalgamated Free Products
In this section we describe the Nielsen reduction method in amalgamated free products. This method was originally developed by H. Zieschang in [29] and then refined by the third author in [20, 21] and, together with R.N. Kalia, in [10].
@@
Next we fix a system $L_{i}$ of left coset representatives of $A$ in $H_{i}$ for $i=1,2$. Here we represent $A$ by $1$. The following proposition is well-known (see [5, Thm. 1.5.8 and Cor. 1.5.10]).
@@
For $g\in G$ with reduced form $g=h_{1}\cdots h_{m}a$, the number $m$ is called the length of $g$. For our purposes, the partial ordering for $G$ given by the length is too coarse. Thus we have to refine it. For $i=1,2$, we take the elements of $L_{i}^{-1}$ as a system of representatives for the right cosets of $A$ in $H_{i}$. Now we have the following normal form (see [5, Sect. 1.5.2] or [28, Prop. 2.3.12]).
@@
    The set $\{g_{1},\dots,g_{n}\}$ is called shorter than the set $\{h_{1},\dots,h_{n}\}$ if $g^{\prime}_{i}\preccurlyeq h^{\prime}_{i}$ for $i=1,\dots,n$ and if $h^{\prime}_{j}\preccurlyeq g^{\prime}_{j}$ fails for at least one $j\in\{1,\dots,n\}$.
@@
Using these rules, the following result was shown by H. Zieschang in [29, Satz 1 and Korollar 1] and refined by the third author in [20, 21].
@@
Finally, we recall some refinements of the preceding material which were shown in [10] and will be useful later. For an element $g\in G$ with symmetric normal form $g=\ell_{1}\cdots\ell_{m}\,k\,r_{m}\cdots r_{1}$ as in Proposition 3, we write $\ell(g)=\ell_{1}\cdots\ell_{m}$ for the leading half, $k(g)=k$ for its kernel, and $r(g)=r_{m}\cdots r_{1}$ for its rear half. Thus we have $g=\ell(g)\cdot k(g)\cdot r(g)$.
@@
Claim (a) is shown in [10, Lemma 2.4] and claim (b) in [10, Lemma 2.5].
@@ 3. Low Rank Subgroups of Amalgamated Free Products
Notice that this notion is completely different from the concept of an “$n$-free group” which usually means that every $n$-generated subgroup is free (see [8]).
In the following we examine subgroups of an amalgamated free product $G=H_{1}\underset{A}{\ast}H_{2}$ which satisfies Assumption 2. In particular, let $H_{1},H_{2}$ be finitely generated groups, and let $A=H_{1}\cap H_{2}$ be an infinite cyclic group which is malnormal in $G$. In this setting, 2-generated subgroups of $G$ can be characterized explicitly. Note that the following result was first shown in [12, Thm. 6] and is included here for completeness sake.
@@
Parts (a) and (b) of this corollary were previously shown in [7]. The next example shows that part (c) is optimal.
@@
Let $G$ be a group of F-type which is not a free product of cyclics. Assume that every subgroup of $G$ of infinite index is a free product of cyclics. Then $G$ is a co-compact planar discontinuous group (cf. [5, Thm. 3.5.45]).
@@ 4. Free Subgroups of Infinite Index
The central open question for free subgroups of infinite index is the Surface Group Conjecture. The first version of this conjecture was formulated 1980 in the Kourovka Notebook by O.V. Melnikov (cf. [14, Problem 7.36]). It has since been disproved, but numerous improved versions were formulated over time (for instance, see [4, 2, 7]), and the latest one reads as follows (cf. [26]).
@@
A celebrated result of H. Wilton is that the surface group conjecture holds for infinite one-relator groups (cf. [26]). Its proof is based on a number of deep results in topology. Another important situation where this conjecture holds is the case of fully residually free groups (see [2]).
@@
For this kind of group, i.e., for $G=\langle H_{1},H_{2};\,u=v\rangle$, where $H_{1}$ and $H_{2}$ are free groups and where $u\in H_{1}$ and $v\in H_{2}$, the surface group conjecture was shown in [2, Thm. 3.1.1]. For the convenience of the reader, we briefly recall the main steps. We distinguish four cases depending on whether $u$ and $v$ are proper powers of elements of $G$.
@@
Case 4: Assume that neither $u$ nor $v$ is a proper power in $G$. In this case, [9, Thm. A] shows that $G$ is a word hyperbolic group. Next we use [24, Thm. 4.11] to show that $G$ is one-ended. In fact, if $G$ has no ends, then $G$ is finite. If $G$ has two ends, then $G$ has an infinite cyclic subgroup of finite index. Finally, if $G$ has more than two ends, it has infinitely many ends and is a non-trivial free product. As all of these cases are impossible for the groups $G$ under consideration, they have to be one-ended. The final step is an application of H. Wilton’s result that a hyperbolic, one-ended cyclically pinched one-relator group, all of whose subgroups of infinite index are free, is a surface group (cf. [27, Corollary 4]).
@@
Not much seems to be known for low rank subgroups and free subgroups of infinite index of HNN extensions. For 2-generated and 3-generated subgroups, the following result from [6] is the best we know. If we drop the assumption that $u$ is not conjugate in $H$ to $v$ or $v^{-1}$, a slightly stronger version for 2-generated subgroups of $G$ is given in [5, Thm. 1.6.42].
@@
As regards to the surface group conjecture, we can show it for the following type of HNN extensions. Notice that the following result is slightly stronger than [2, Thm. 3.1.2] and [5, Thm. 1.7.22].
@@
Let $N$ be the normal closure of $H$ in $G$. Then $G/N$ is an infinite cyclic group. Hence the group $H$ is of infinite index in $G$, and the hypothesis implies that $H$ is a free group. Thus $G$ is a conjugacy pinched one-relator group. For such groups, the surface group conjecture can be shown based in part on [2, Thm. 3.1.2]. Let us follow the main steps and complete the proof at the appropriate place. We distinguish two cases depending on whether $u$ or $v$ are proper powers of elements of $G$.
Case 1: Assume that $u=g^{m}$ and $v=h^{n}$ with $g,h\in G$ and with $m,n\geq 2$. If $G=\langle g,h\mid g^{2}=h^{2}\rangle$ then $G$ is a surface group and contains a free abelian subgroup of rank 2. Now assume that $G$ is not isomorphic to $\langle g,h\mid g^{2}=h^{2}\rangle$. Using the normal form in HNN groups (cf. [5, Thm. 1.6.4]), it follows that the subgroup generated by $w=tgt^{-1}$ and $h$ has the presentation $W=\langle w,h\mid w^{m}=h^{n}\rangle$. Then the subgroup $\langle w^{m},wh\rangle$ of $W$ is a free abelian group of rank 2 and has infinite index in $G$, in contradiction to the hypothesis.
@@
In the case $\mathop{\rm rk}\nolimits(H)=1$, we let $H=\langle a\rangle$. Then we get $G=\langle a,t\mid tat^{-1}=a^{k}\rangle=B(1,k)$ with a Baumslag-Solitar group $B(1,k)$. For $|k|=1$, this is a surface group, and for $p=|k|\geq 2$, the group $B(1,p)$ is isomorphic to $\mathbb{Z}[\frac{1}{p}]\rtimes\mathbb{Z}$ (cf. [3]) and contains $\mathbb{Z}[\frac{1}{p}]$ as a non-free subgroup of infinite index. Hence, also in the case $G\cong B(1,k)$ with $|k|\geq 2$, the group $G$ has a non-free subgroup of infinite index, in contradiction to the hypothesis.
Subcase 2b: Assume that $\langle u\rangle\cap x\langle v\rangle x^{-1}$ is finite for every $x\in H$. Since $G$ is torsion free, we get $\langle u\rangle\cap x\langle v\rangle x^{-1}=\{1\}$. By [13], it follows that $G$ is word hyperbolic. As in the proof of Theorem 16, we see that $G$ is one-ended, and then an application of [27] finishes the proof.